OpenStudy (laronjaxon):

Solving Quadratics by Factoring PLEASE 2. Solve x2 = 14x – 49

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

x2 - 14x + 49 = 0 x2 - 7x - 7x + 49 = 0 Taking common, x(x - 7) -7(x - 7) = 0 So, (x - 7)(x - 7) = 0 Therefore, x = 7 is the answer..

OpenStudy (laronjaxon):

THANKS (:

OpenStudy (anonymous):

Welcome dear..